OpenStudy (anonymous):

Michelle uses a credit card with a 23.2% APR compounded monthly to pay for car repairs totaling $790.59. She can pay $330 per month on the card. What will the total cost of this purchase be? $974.01 $790.59 $818.07 $948.66

OpenStudy (anonymous):

Its going to be 818.07 because first month interest is ~15 dollars second month interest is ~ 9 dollars final month interest will be about ~ 2

OpenStudy (anonymous):

take the initial value multiple it by the interest rate divide and you have your interest take the interest amount and divide it by 12 to get your monthly interest then add that to you initial value

OpenStudy (anonymous):

from there subtract your 330 and repeat that process till the last month

OpenStudy (anonymous):

So it would be 790.59(1+0.232)^12-330

OpenStudy (anonymous):

close forgive me for not remembering the exact formula you need to use but you go like this 790.59*.232 = 183.41688 / 12 = 15.28474 + 790.59 = 805.87474 as your first month balance then subtract the 330 and use your initial to start the numbers again you don't worry about the exponent because you are just finding the interest for the single month. Your total balance changes each month and the interest charged will change with it.

OpenStudy (anonymous):

you're actually close 790.59(1 + .232/12) ^1 - 330

OpenStudy (anonymous):

So with that I got 475.87, what do you do after that?

OpenStudy (anonymous):

take that number and put it into the formula again 475.87474(1+.232/12)^1 -330 = 155.074985 155.074985(1.232/12^1 = 158.0731013 is your last payment due so add up 330 + 330 + 158.0731013 = 818.0731013

OpenStudy (anonymous):

Ooooooh okay, thank you!

OpenStudy (anonymous):

Would I use that same equation for this problem too?? Ken owes $6,451 on a credit card with an 18.6% interest rate compounded monthly. What is the monthly payment he should make in order to pay off this debt in 12 months, assuming he does not charge any more purchases with the card? $112.24 $593.27 $1,377.78 $537.58

OpenStudy (anonymous):

its a little more difficult this time. I wish I remembered the exact equation because it would make your life so much easier. I can tell you the answer will be either B or D. Give me one second and I will try and type this all out.

OpenStudy (anonymous):

Okay, thank you though! I'm thinking it's B

OpenStudy (anonymous):

its definitely B because D is what you would pay monthly with no interest:)

OpenStudy (anonymous):

I need to leave work and havent been able to figure it out, if you havent gotten an answer to this later tonight I will try and give it another go. B makes sense but I can't get the number to match up for some reason.

OpenStudy (anonymous):

Okay, thank you. I'll tell you if the answer is correct or not when I get it